There are two types of individuals in the demand for natural gas. Group X has a total of 5 people and each person’s demand curve is Q = 100 – 10*P while Group Y has 10 people and each person’s demand curve is Q = 100 – 5*P. a) Assume natural gas costs $10 per unit.

a)Calculate the consumer surplus amounts for each person (for a person in Group X and a person in Group Y).

b) Construct the aggregate demand curve for Group X. Also construct the aggregate demand curve for group Y.

c) Construct the market demand curve for this economy (all people). Do this graphically or mathematically.
Answer #1

Please find the solution attached below

for Group 6,= 100 - SP 100 - By [ Indiuided demand conte] P PMC-10 TUU- & 10 > y = 50 umsame Suplus 21 x bxka = *sox (90-10)(6) For Group x Aggregate demand for Group is given nx [100-10P] = 500-50P - 0 For Group Y na 10 n 100 -SP] 0 10 [100 -SP] 10